• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name weinstein

    Log entry time 17:12:52 on October29,2001

    Entry number 75407


    espace replay of run 1364 detectors found some problems:
    1) hadron s1 and s2 ADC are badly gain matched
    2) there are peculiar dips in the left s0 tdc spectrum and it has a double peak
    3) left arm VDC plane 2 has about 0.5 fewer hits per cluster than the other planes although the v1 coarse spectrum looks fine